OpenStudy (anonymous):

Medal for help. What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (shiraz14):

-2

OpenStudy (mathmale):

Have you seen and used the slope formula before? if so, please type it out here.

OpenStudy (yanasidlinskiy):

m=y2-y1/x2-x1

OpenStudy (anonymous):

Thanks guys

OpenStudy (shiraz14):

The answer's just for checking purpose - you should know how to use the formula (as @mathmale has stated)

OpenStudy (anonymous):

use slope formula: \[m= slope\] so, m= y (sub 2) - y (sub 1) / x (sub 2) - x (sub 1) now let's substitute: m= 3-(7)/ -1-4 then simplify.....

OpenStudy (shiraz14):

@FOXSHADOW : Slight typo in your post above - it should be 3-(-7), not 3-(7) ...

OpenStudy (anonymous):

@shiraz14 ha sorry I completely missed that... thanks

OpenStudy (shiraz14):

@FOXSHADOW : You're welcome